OpenStudy (anonymous):

what is 3 times 7 squared by 7?

OpenStudy (benhogi):

i think its 3 but i have no idea

OpenStudy (anonymous):

thanks

OpenStudy (er.mohd.amir):

3*sq 7/7 =3*7*7/7 =21

OpenStudy (er.mohd.amir):

this is right